# VRWKV-Editor: Reducing Quadratic Complexity in Transformer-Based Video Editing

### v2

@@ 5 Experiments
We conduct experiments to test the performance of VRWKV-Editor with existing video-editing methods. Our evaluation encompasses computational efficiency metrics, generation quality assessments, and temporal coherence analysis across various video editing tasks.
5.1 Implementation Details
During inference, we implemented the DDIM sampler [47] alongside classifier-free guidance [48] of $7.5$, controlling how strongly the model follows the text prompt. All experiments are conducted on a single machine with an NVIDIA A100-SXM4-80GB GPU.
5.2 Computational Efficiency Analysis:
Timing Performance Comparison: Our timing experiments reveal significant computational advantages of VRWKV-Editor over window-based attention architectures. Figure 6 presents the forward and backward pass timing results for different input sequence lengths. The results demonstrate that VRWKV-Editor achieves substantial speedups that increase with input size, confirming the theoretical $O(n)$ vs $O(n^{2})$ complexity advantage. For large inputs (4096×4096), our method achieves up to 3.68× speedup in total computation time.
Memory Efficiency Analysis: VRWKV-Editor exhibits significantly better memory scaling compared to traditional attention-based architectures. While memory usage of conventional attention mechanisms grows quadratically with sequence length due to attention matrix storage, VRWKV-Editor maintains linear memory growth. At 4K resolution, VRWKV-Editor uses approximately 60% less GPU memory than swin attention based approaches.
Quantitative Results: In Table 1, we report the quantitative results of VRWKV-Editor versus state-of-the-art video editing methods. As the results show in Figures 1, 3 and Table 1, our model achieves a trade-off between temporal consistency, text alignment (for objects, background, and style) and high-quality video editing while significantly reducing computational complexity. This is also evidenced by the performance obtained in both automatic evaluations and user studies.
Qualitative Evaluation: Figure 3 showcases a visual comparison between our method and the other baselines. Further results are shown in Figures 7 and 8. We conducted our evaluation by editing objects, background, style or combination between them.
Table 1: Quantitative comparison between our method and evaluated baselines in terms of frame consistency, textual alignment, and runtime (minutes).

### v3

@@ 5 Experiments
We conduct experiments to test the performance of VRWKV-Editor with existing video editing methods. Our evaluation encompasses computational efficiency metrics, generation quality assessments, and temporal coherence analysis across various video editing tasks.
Table 1: Quantitative comparison between our method and evaluated baselines in terms of frame consistency, textual alignment, and runtime for overall pipelines (minutes).
@@
5.1 Implementation Details
We used the pretrained weights of Stable Diffusion v1.5 [40] to initialize the 2D models of our framework. In the training stage, the AdamW optimizer is employed with a learning rate of $3e^{-5}$ for a total of 300 to 500 steps with 100 validation steps and a batch size of one video. For each video, we designed 3 edited prompts covering objects, background, color, weather, or style changing. During inference, we implemented the DDIM sampler [52] alongside classifier-free guidance [53] of $7.5$, controlling how strongly the model follows the text prompt. All experiments are conducted on a single machine with an NVIDIA A100-SXM4-80GB GPU.
@@ 5.1 Implementation Details
Dataset and Metrics: To evaluate our approach, we extended the Video-P2P dataset [14] by incorporating 13 additional videos from the DAVIS dataset [54], resulting in a total of 24 videos. Each video was paired with three text prompts, yielding 72 edited samples overall. For evaluation, we report CLIP-Score and Pick-Score to measure semantic alignment and perceptual quality, respectively, alongside user preference studies and runtime analysis.
5.2 Computational Efficiency Analysis:
Timing Performance Comparison: Our timing experiments reveal significant computational advantages of VRWKV-Editor over window-based attention architectures. Figures 6 and 7 presents the forward and backward pass timing results for different input sequence lengths, and running time of end-to-end editing pipelines for all evaluated methods are reported in Table 1. The results demonstrate that VRWKV-Editor achieves substantial speedups that increase with input size, confirming the theoretical $O(n)$ vs $O(n^{2})$ complexity advantage. For large inputs (4096×4096), our method achieves up to 3.68× speedup in total computation time.
Quantitative Results: In Table 1, we report the quantitative results of VRWKV-Editor versus state-of-the-art video editing methods. As the results shows, our model achieves a trade-off between temporal consistency, text alignment (for objects, background, and style) and high-quality video editing while significantly reducing computational complexity. This is also evidenced by the performance obtained in both automatic evaluations and user studies.
Qualitative Evaluation: Figure 2 showcases a visual comparison between our method and the other baselines. Further results are shown in Figure 1 and on the project page VRWKV-Editor. We conducted our evaluation by editing objects, background, style or combination between them.
@@ 6 Limitations:
Finally, our evaluation relies primarily on CLIP-based scores and user studies, which, while useful, do not fully capture video-specific qualities such as temporal consistency and motion smoothness. Future work should therefore prioritize adapting VRWKV-Editor to its RNN mode for scalable long-sequence editing, and adopting more sophisticated video-centric evaluation metrics.